# Preempting Fermion Sign Problem: Unveiling Quantum Criticality through Nonequilibrium Dynamics

### v1

@@ Abstract
The notorious fermion sign problem, arising from fermion statistics, constitutes one of the main obstacles of deciphering quantum many-body systems by numerical approach. The progress in overcoming sign problem will definitely lead to a great leap in various areas of modern physics. Here, by deviating from the conventional cognition that nonequilibrium studies should be more complicated than equilibrium cases, we propose an innovative framework based on nonequilibrium critical dynamics to preempt sign problem and investigate quantum critical point in fermionic model through numerically exact quantum Monte Carlo (QMC) simulation. By virtue of universal scaling theory of imaginary-time relaxation dynamics, we demonstrate that accurate critical point and critical exponents can be obtained in the short-time stage, in which the sign problem is not severe such that the QMC is accessible. After confirming the effectiveness of the method in two typical interacting fermionic models featuring Dirac quantum critical point (QCP), we for the first time reveal the quantum phase diagram in the Hubbard model hosting $\rm SU(3)$-symmetric Dirac fermions, and find that the QCP between Dirac semi-metal and a $\lambda_{8}$-antiferromagnetic phase belongs to a new universality class different from the previously known Gross-Neveu transitions.
11footnotetext: These authors contribute equally to this work.
@@
Introduction — The sign structure of the quantum many-body wave functions works like a double-edged sword. On the one hand, it endows fertile exotic phenomena ranging from condensed matter physics [1, 2] to high energy physics [3]. Quantum Monte-Carlo (QMC) is among the most important theoretical approaches to study the exotic properties of strongly correlated quantum many-body systems [4, 5, 6], because it is intrinsically unbiased and numerically exact. However, on the other hand, the sign structure in the wave function results in the notorious sign problem [7, 8, 9, 10, 11], which considerably plagues the application of QMC to investigating interacting models potentially featuring intriguing physics, for instance, the Hubbard model at generic filling [12, 13] and the lattice QCD at finite baryon density [14]. Since the general solution of sign problem is lacking and nondeterministic polynomial (NP)-hard [9], establishing novel generic strategies to attack the sign problem in QMC will definitely provoke significant advances in the realm of quantum many-body physics [15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36].
@@
Among various exotic phenomena in quantum many-body systems, quantum criticality emerges as a particularly crucial and fascinating one. The underlying physics of the quantum critical point (QCP) lays the foundation for achieving a unified theoretical framework to characterize different phases and phase transitions [37, 38, 39, 40, 41]. Moreover, quantum criticality is intimately associated with fundamentally important phenomena in condensed matter physics such as high-$T_{c}$ superconductivity [42, 43, 44] and strange metallicity [45, 46]. The understanding of QCP in a non-perturbative way is severely hindered by the sign problem likewise. Developing a generic unbiased theoretical approach to decipher quantum criticality not only has overarching meaning in fundamental theory, but also immensely promotes the research on the exotic phenomena emerging in quantum materials.
@@
Figure 1: Scheme of preempting sign problem to probe quantum criticality via short-time critical dynamics and the application in single-Dirac-fermion Hubbard model. a, With some typical initial states, scaling behaviors governed by the QCP are reflected in the short-time stage, in which the sign problem is still weak as shown in the inset. The red dashed line therein marks the average sign at $\tau=0.3L^{z}$. b, Determination of QCP as $U_{c}=7.220(37)$ via the intersection points of curves of the correlation length ratio $R_{\rm FM}$ versus $U$ for different sizes at $\tau=0.3L^{z}$ with DSM initial state. c, Determination of $1/\nu=1.18(3)$ by scaling collapse of $R_{\rm FM}$ versus rescaled $(U-U_{c})$. d, Determination of $\eta_{\phi}=0.33(2)$ via scaling collapse of curves of the structure factor $S_{\rm FM}$ versus rescaled $\tau$ at $U_{c}$. e, Determination of $\eta_{\psi}=0.135(2)$ via scaling collapse of curves of the fermion correlation $G_{\rm f}$ versus rescaled $\tau$ at $U_{c}$.
@@
Here, we propose a general framework to preempt sign problem in QMC and unveil the quantum criticality. Generally, fathoming the nonequilibrium properties in quantum many-body systems is substantially challenging, compared with the equilibrium ones. However, we leverage the nonequilibrium behavior and demonstrate that the short-time critical dynamics, which was first proposed in classical systems [47] and used to detect the classical critical behaviors [48] and then generalized to imaginary-time dynamics of QCP [49], provides an ingenious strategy to reliably probe quantum criticality in the presence of sign problem, as illustrated in Fig. 1a. The underling mechanism is that in the short-time stage of imaginary-time relaxation for some typical initial states, (1) universal scaling behaviors manifesting the quantum criticality appear  [49, 50]; (2) the sign problem remains mild in this stage such that the reliable results are accessible by QMC simulation. Hence, bestowing the scaling of short-time dynamics [47, 48, 49, 50], we can accurately determine critical properties of the QCP, while the sign problem therein, whose severity generally exponentially increases with the imaginary time [9, 51], is largely alleviated compared with the equilibrium one involving long imaginary-time evolution.
@@
In the following, we demonstrate the framework by studying two typical interacting fermionic models featuring Dirac QCP. We unambiguously show that accurate critical properties are accessed through short-time dynamics, despite the presence of sign problem. More remarkably, we adopt the approach to investigate a paradigmatic strongly interacting model hosting Dirac fermions with $\rm SU(3)$ symmetry, namely the $\rm SU(3)$ staggered-flux Hubbard model, which is sign problematic in any known algorithm. For the first time we unravel the ground-state phase diagram of the model and reveal the critical properties of the quantum phase transition between the Dirac semi-metal (DSM) and a $\lambda_{8}$-antiferromagnetic (AFM) phase. Intriguingly, the QCP belongs to a novel universality class distinct from previous Gross-Neveu transitions.
@@
Theoretical framework — We consider the imaginary-time relaxation dynamics for which the wave function $|\psi(\tau)\rangle$ evolves according to the imaginary-time Schrödinger equation $-\frac{\partial}{\partial\tau}|\psi(\tau)\rangle=H|\psi(\tau)\rangle$ imposed by the normalization condition. As its long-time solution is the ground state, this equation provides a routine method to access ground-state properties in numerical computations [51]. Moreover, near the QCP, it was shown that universal scaling behaviors manifest themselves not only at the ground state, but also in the imaginary-time relaxation process. Particularly, with initial states corresponding to the fixed points of scale transformation, the short-time dynamics of observable $O$ satisfies the following scaling form [49, 50]:
@@
where $L$ is linear system size and $g$ is the distance to the critical point, $\kappa$ is the scaling dimension of $O$, and $\nu$ is the correlation-length exponent, $z$ is dynamical exponent which is generally one in Dirac QCP. Note that all critical exponents in Eq. (1) are controlled by the QCP in the ground state. Accordingly, from Eq. (1), one can achieve critical properties from the expectation value of $O$ in the relaxation process. This nonequilibrium approach is notably different from the conventional one which requires that the ground state should be determined through the evolution of long imaginary time. Because in general the severity of sign problem exponentially increases with $\tau$ in QMC [9], the sign problem in short-time stage of relaxation is significantly mitigated compared with the simulation on ground-state properties, thereby enabling the large-scale QMC simulation with high accuracy despite the presence of sign problem. In the following, we elucidate the theoretical framework by systematically studying several representative models with sign problem.
@@
Single-Dirac-fermion Hubbard model — We first consider the SLAC fermion Hubbard model with the Hamiltonian [52, 53, 54]:
$H=\sum_{\bm{i},\bm{R}}t_{\bm{R}}c^{\dagger}_{\bm{i}\uparrow}c_{\bm{i}+\bm{R}\downarrow}+\mathrm{h.c.}+U\sum_{i}\left(n_{i\uparrow}-\frac{1}{2}\right)\left(n_{i\downarrow}-\frac{1}{2}\right),$
@@
where $t_{\bm{R}}=\frac{i(-1)^{R_{x}}}{\frac{L}{\pi}\sin{\pi\frac{R_{x}}{L}}}\delta_{R_{y},0}+\frac{(-1)^{R_{y}}}{\frac{L}{\pi}\sin{\pi\frac{R_{y}}{L}}}\delta_{R_{x},0}$ is the amplitude of long-range hopping and $U$ is the strength of repulsive interaction. The model is sign problematic in any known QMC algorithm. Nonetheless, since the sign problem is relatively benign, previous QMC study reveals a chiral Ising QCP separating Dirac semimetal (DSM) phase and ferromagnetic (FM) phase [52].
@@
Figure 2: Probing quantum criticality via short-time critical dynamics in spinless $t$-$V$ model with CDW initial state. a, Determination of QCP as $V_{c}=1.35(1)$ via the intersection points of $R_{\rm CDW}$ versus $V$ for different $L$ at $\tau=0.3L^{z}$. Shown in Inset is the evolution of average sign with red dashed curve marks $\tau=0.3L^{z}$. b, Determination of $1/\nu=1.30(18)$ via scaling collapse of $R_{\rm CDW}$ versus rescaled $(V-V_{c})$. c-d, Determination of $\eta_{\phi}=0.49(5)$ and $\eta_{\psi}=0.073(4)$ via scaling collapse of short-time dynamics of $S_{\rm CDW}$ and $G_{\rm f}$ versus rescaled $\tau$, respectively.
@@
Here, we present the procedure for unraveling the critical properties via the short-time scaling of the observable as dictated in Eq. (1). In the conventional projector QMC approach, a sufficiently long evolution time should be implemented to ensure that the ground-state properties are accessed, resulting in the severe sign problem exponentially increasing with $\tau$, as shown in Fig. 1. We overcome this critical difficulty owing to the short-time critical relaxation dynamics. To determine the critical point via Eq. (1), we consider observable $O$ as the dimensionless correlation length ratio $R_{\rm FM}$ for the FM order (See SM Sec. II for the detailed definition [55]) and fix $\tau$ at short-time stage as $\tau=0.3L^{z}$, where $z=1$ due to the nature of Dirac QCP. Accordingly, $R_{\rm FM}$ satisfies $R_{\rm FM}=f_{R}(gL^{1/\nu})$ where $g=U-U_{c}$. We show the numerical results in Fig. 1b, in which the crossing point of $R_{\rm FM}$ versus $U$ for different $L$ reveals a quantum phase transition from DSM to FM phase occurring at $U_{c}=7.220(37)$. Furthermore, we perform data collapse analysis of $R_{\rm FM}$ for different $L$ versus $(U-U_{c})L^{1/\nu}$ and achieve accurate critical exponent $1/\nu=1.18(3)$, as shown in Fig. 1c. Furthermore, fixing $U=U_{c}$, we calculate the imaginary-time evolution of FM structure order $S_{\rm FM}$ and fermion correlation $G_{\rm f}$ (defined in SM Sec. II [55]), whose scaling dimensions are $(1+\eta_{\phi})$ and $\eta_{\psi}$, respectively [52, 50]. By making the rescaled curves of $S_{\rm FM}L^{(1+\eta_{\phi})}$ and $G_{\rm f}L^{\eta_{\psi}}$ versus $\tau L^{-z}$ for different $L$ collapse according to Eq. (1) as shown in Figs. 1d and 1e, we achieve the anomalous dimensions of the order parameter boson $\eta_{\phi}=0.33(2)$ and the Dirac fermion $\eta_{\psi}=0.135(2)$, respectively.
Comparing the critical point and critical exponents evaluated here with previous results, we find that all are approximately consistent with previous study of Gutzwiller QMC [52] and functional renormalization group  [56]. Notice that since sign problem is largely alleviated in our work owing to the short evolution time, we can achieve larger system size $L=23$ [55], compared with previous studies, and thus reach the reliable results of critical exponents. Consequently, we unambiguously demonstrate that the short-time dynamics via QMC simulation provides a powerful approach to fathom quantum critical properties in sign-problematic strongly interacting models.
@@
Spinless $t$-$V$ model — To further demonstrate the framework of our approach, we study another typical interacting Dirac-fermion model, termed as Honeycomb spinless $t$-$V$ model, with the Hamiltonian:  [17, 57, 58, 59]:
@@
where $t$ is nearest-neighbor (NN) hopping amplitude and $V$ denotes the strength of NN density interaction. The appearance of sign problem in Eq. (3) depends on the channel of Hubbard-Strotonovich (H-S) transformation. Hence, the model provides a genuine platform to confirm the accuracy and feasibility of our approach to unravel QCP in the presence of sign problem. Previous sign-free QMC studies, with the H-S transformation in the hopping channel, reveal the critical properties of the QCP separating DSM and charge-density-wave (CDW) phases [58]. Here, we decouple the interaction in the sign-problematic density channel and study the QCP via short-time relaxation dynamics [60].
@@
Figure 3: Phase diagram and quantum criticality in $\rm SU(3)$ Hubbard model detected by short-time dynamics with $\lambda_{8}$-AFM initial state. a, Phase diagram determined via short-time dynamics. Insets show the energy spectra of DSM state (upper left) and Sketch of $\lambda_{8}$-AFM order in which a fermion with one flavor (blue) is situated at one sublattice and double fermions with the other two flavors (pink and violet) are situated at the other sublattice (lower right). b-c, Critical point $U_{c}=1.10(5)$ for $\phi=0.075\pi$ and $1/\nu=0.68(5)$ determined via curves of $R_{\lambda_{8}\text{-AFM}}$ versus $U$ for different $L$ at $\tau=0.25L^{z}$. Shown in Inset of b is the evolution of average sign with red dashed curve marks $\tau=0.25L^{z}$. d-e, $\eta_{\phi}=0.55(5)$ and $\eta_{\psi}=0.15(3)$ determined via the scaling collapse of evolution curves of $S_{\lambda_{8}\text{-AFM}}$ and $G_{\rm f}$, respectively.
@@
The procedure is the same as the previous section, except that the initial state is chosen as the CDW fully ordered state. We will demonstrate that the initial state can be flexibly selected in our approach, thereby offering various routes to extract critical properties. We fix $\tau=0.3L^{z}$, and determine QCP from the crossing point of curves of correlation-length ratio $R_{\rm CDW}$ versus $V$ ($R_{\rm CDW}$ is defined in SM Sec. III [55]), as displayed in Fig. 2a, giving rise to $V_{c}=1.35(1)$. The scaling collapse analysis determines the value of $\nu$ as $1/\nu=1.30(18)$. Additionally, the anomalous dimensions $\eta_{\phi}=0.49(5)$ and $\eta_{\psi}=0.073(4)$ are obtained from the evolution of the CDW structure factor $S_{\rm CDW}$ and the fermion correlation $G_{\rm f}$ (defined in SM Sec. III [55]), respectively, at $V_{c}$. Both $V_{c}$ and $\eta_{\phi}$ are consistent with previous numerical results, while $\eta_{\psi}$, to the best of our knowledge, is numerically determined for the first time, and is consistent with the previous results of the functional renormalization group [56]. The similar computation with DSM initial state yields the consistent results, with the details included in the SM [55]. Hence, these results for model (3) further establish the feasibility of the approach based on short-time dynamics in accessing the critical properties in the presence of sign problem. Moreover, the flexible choice of initial states provide benchmarks for the results, further improving the accuracy of the approach.
@@
SU(3) Hubbard model — The successful application of the method in previous two models encourages us to study the critical properties in unexplored models. In the following, we implement the approach to systematically investigate a sign-problematic interacting model featuring exotic Dirac QCP belonging to a novel universality class. We consider $\rm SU(3)$ Hubbard model with staggered magnetic flux on square lattice, described by the following Hamiltonian:

### v2

@@ Abstract
The notorious fermion sign problem, arising from fermion statistics, presents a fundamental obstacle to the numerical simulation of quantum many-body systems. Here, we introduce a framework that circumvents the sign problem in the studies of quantum criticality and its associated phases by leveraging imaginary-time nonequilibrium critical dynamics. We demonstrate that the critical properties can be accurately determined from the system’s short-time relaxation, a regime where the sign problem remains manageable for quantum Monte-Carlo (QMC) simulations. After validating this approach on two benchmark fermionic models, we apply it to the sign-problematic Hubbard model hosting SU(3)-symmetric Dirac fermions. We present the first numerically exact characterization of its quantum phase diagram, revealing a continuous transition between a Dirac semi-metal and a SU(3) antiferromagnetic phase. This transition defines an unconventional Gross-Neveu universality class that fundamentally reshapes current understanding of Gross-Neveu criticality. Our work provides a powerful tool for investigating sign-problematic systems and quantum criticality.
INTRODUCTION
@@
The complex sign structure of the quantum mechanical wavefunctions presents a central dichotomy in many-body physics. On the one hand, it endows fertile exotic phenomena across condensed matter [1, 2] and high energy physics [3]. Quantum Monte-Carlo (QMC) is among the most important theoretical approaches to study the strongly correlated quantum many-body systems [4, 5, 6], offering numerically exact and unbiased solutions. On the other hand, the sign structure results in the notorious sign problem [7, 8, 9, 10, 11], which considerably plagues the application of QMC to investigating interacting models potentially featuring intriguing physics, including the Hubbard model at generic filling [12, 13] and the lattice QCD at finite baryon density [14]. Since the general solution of the sign problem is lacking and nondeterministic polynomial (NP)-hard [9], developing generic strategies to circumvent or mitigate it will definitely provoke significant advances in the realm of quantum many-body physics [15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36].
@@
Among various exotic phenomena in quantum many-body systems, quantum criticality emerges as a particularly crucial and fascinating one. The underlying physics of the quantum critical point (QCP) lays the foundation for achieving a unified theoretical framework to characterize different phases and phase transitions [37, 38, 39, 40, 41]. Moreover, quantum criticality is intimately associated with fundamentally important phenomena in condensed matter physics such as high-$T_{c}$ superconductivity [42, 43, 44] and strange metallicity [45, 46]. The understanding of the QCP in a non-perturbative way is severely hindered by the sign problem likewise.
@@
Figure 1: Scheme of preempting sign problem to probe quantum criticality via short-time critical dynamics and the application in single-Dirac-fermion Hubbard model. a, With some typical initial states, scaling behaviors governed by the QCP are reflected in the short-time stage, in which the sign problem is still weak as shown in the inset. The white dashed line indicates the nonequilibrium critical region associated with the QCP. The red dashed line in the inset indicates the average sign at $\tau=0.3L^{z}$, where $z=1$ due to the nature of the Dirac QCP. b, Determination of QCP as $U_{c}=7.220(37)$ via the intersection points of curves of the correlation length ratio $R_{\rm FM}$ versus $U$ for different sizes at $\tau=0.3L^{z}$ with DSM initial state. c, Determination of $1/\nu=1.18(3)$ by scaling collapse of $R_{\rm FM}$ versus rescaled $(U-U_{c})$, with a reduced chi-squared value $\chi_{\nu}^{2}=1.824$ indicating good quality. d, Determination of $\eta_{\phi}=0.33(2)$ via scaling collapse of curves of the structure factor $S_{\rm FM}$ versus rescaled $\tau$ at $U_{c}$, with $\chi_{\nu}^{2}=0.606$. e, Determination of $\eta_{\psi}=0.135(2)$ via scaling collapse of curves of the fermion correlation $G_{\rm f}$ versus rescaled $\tau$ at $U_{c}$, with $\chi_{\nu}^{2}=1.241$. The dashed lines in d and e indicate the boundary of the nonequilibrium scaling region; only the regions to the right are included in the scaling collapse analysis (see SM Sec. S2.2).
@@
Here, we propose a general framework to preempt the sign problem in QMC and unveil the quantum criticality and associated ground-state phase diagram. It is widely acknowledged that fathoming the nonequilibrium properties in quantum many-body systems is substantially more challenging than equilibrium ones. However, we leverage the nonequilibrium behavior and demonstrate that the short-time critical dynamics, which was first proposed in classical systems [47] and used to detect the classical critical behaviors [48] and then generalized to imaginary-time dynamics of QCP [49], provides an ingenious strategy to reliably probe quantum criticality in the presence of the sign problem, as illustrated in Fig. 1a. The underlying mechanism is that in the short-time stage of imaginary-time relaxation for some typical initial states, (1) universal scaling behaviors manifesting the quantum criticality appear [49, 50]; (2) the sign problem remains mild in this stage, such that the reliable results are accessible by QMC simulation. Hence, bestowing the scaling of short-time dynamics [47, 48, 49, 50], we can accurately determine the location and critical properties of the QCP, while the sign problem therein, whose severity generally increases exponentially with the imaginary time [9, 51, 52], is largely alleviated compared with the equilibrium one involving long imaginary-time evolution.
@@
In the following, we demonstrate the framework by studying two typical interacting fermionic models featuring Dirac QCP. We unambiguously show that accurate critical properties can be accessed through short-imaginary-time dynamics, despite the presence of the sign problem. More remarkably, we adopt the approach to investigate a paradigmatic strongly interacting model hosting Dirac fermions with $\rm SU(3)$ symmetry, which is sign-problematic in any known algorithm. For the first time, we establish the ground-state phase diagram of the model and reveal the critical properties of the quantum phase transition between the Dirac semi-metal (DSM) and a $\lambda_{8}$-antiferromagnetic (AFM) phase. Intriguingly, the QCP defines an unconventional Gross-Neveu transition that goes beyond the previous paradigms of O($N$) ordering, distinct from the known universality classes of Gross-Neveu transition.
RESULTS
Theoretical framework
@@ RESULTS
We consider the imaginary-time relaxation dynamics for which the wave function $|\psi(\tau)\rangle$ evolves according to the imaginary-time Schrödinger equation $\frac{\partial}{\partial\tau}|\psi(\tau)\rangle=H|\psi(\tau)\rangle$ imposed by the normalization condition. As its long-time solution is the ground state, this equation forms the basis of the zero-temperature projector QMC (PQMC) method, which provides a routine approach to accessing ground-state properties in numerical computations [53, 54, 55, 51]. Moreover, near the QCP, it was shown that universal scaling behaviors manifest themselves not only at the ground state, but also in the imaginary-time relaxation process. Particularly, with initial states corresponding to the fixed points of scale transformation, after a transient nonuniversal time scale, the short-imaginary-time dynamics of observable $O$ satisfies the following scaling form [49, 50]:
@@
where $L$ is linear system size and $g$ is the distance to the critical point, $\kappa$ is the scaling dimension of $O$, and $\nu$ is the correlation-length exponent, $z$ is the dynamical exponent which is generally one in the Dirac QCP. Note that all critical exponents in Eq. (1) are controlled by the QCP in the ground state. Accordingly, from Eq. (1), one can achieve the location of QCP and critical properties from the expectation value of $O$ in the relaxation process. This nonequilibrium approach is notably different from the conventional PQMC which requires that the ground state should be determined through the evolution of long imaginary time [53, 54, 55, 51]. Because in general the severity of the sign problem exponentially increases with imaginary time in QMC [9, 52], the sign problem in the short-time stage of relaxation is significantly mitigated compared with the direct simulation on ground-state properties, thereby enabling the large-scale QMC simulation with high accuracy despite the presence of the sign problem. In the following, we elucidate the theoretical framework by systematically studying several representative models with the sign problem.
Single-Dirac-fermion Hubbard model
@@
We first consider the SLAC fermion Hubbard model with the Hamiltonian [56, 57, 58]:
$H=\sum_{i,R}t_{R}c^{\dagger}_{i\uparrow}c_{i+R\downarrow}+\mathrm{h.c.}+U\sum_{i}\left(n_{i\uparrow}-\frac{1}{2}\right)\left(n_{i\downarrow}-\frac{1}{2}\right),$
@@
where $t_{R}=\frac{i(-1)^{R_{x}}}{\frac{L}{\pi}\sin{\pi\frac{R_{x}}{L}}}\delta_{R_{y},0}+\frac{(-1)^{R_{y}}}{\frac{L}{\pi}\sin{\pi\frac{R_{y}}{L}}}\delta_{R_{x},0}$ is the amplitude of long-range hopping on the square lattice and $U$ is the strength of repulsive interaction. The model is sign-problematic in any known QMC algorithm. Nonetheless, since the sign problem is relatively benign, a previous QMC study reveals a chiral Ising QCP separating the DSM phase and ferromagnetic (FM) phase [56].
@@
Figure 2: Probing quantum criticality via short-imaginary-time critical dynamics in spinless $t$-$V$ model with CDW initial state. a, Determination of QCP as $V_{c}=1.35(1)$ via the intersection points of $R_{\rm CDW}$ versus $V$ for different $L$ at $\tau=0.3L^{z}$, where $z=1$. Shown in Inset is the evolution of average sign with red dashed curve marks $\tau=0.3L^{z}$. b, Determination of $1/\nu=1.30(18)$ via scaling collapse of $R_{\rm CDW}$ versus rescaled $(V-V_{c})$, with $\chi_{\nu}^{2}=1.518$ demonstrating the collapse quality. c-d, Determination of $\eta_{\phi}=0.49(5)$ and $\eta_{\psi}=0.073(4)$ via scaling collapse of the short-imaginary-time dynamics of $S_{\rm CDW}$ and $G_{\rm f}$ versus rescaled $\tau$, respectively, with $\chi_{\nu}^{2}=1.668$ and $0.963$ demonstrating the collapse quality in the nonequilibrium scaling region to the right of the dashed lines.
@@
Here, we present the procedure for unraveling the critical properties via the short-imaginary-time scaling of the observable as dictated in Eq. (1). In conventional PQMC, long imaginary-time evolution is typically required [53, 54, 55, 51], resulting in the severe sign problem exponentially increasing with $\tau$, as shown in the inset of Fig. 1a. We overcome this critical difficulty by leveraging short-imaginary-time critical relaxation dynamics, with the DSM state chosen as the initial state for the following demonstration. To determine the critical point via Eq. (1), we consider observable $O$ as the dimensionless correlation length ratio $R_{\rm FM}$ for the FM order (See SM Sec. S2 for the detailed definition [53]). We take a short imaginary-time length $\tau=0.3L^{z}$, where Eq. (1) reduces to $R_{\rm FM}=f_{R}(gL^{1/\nu})$. Here the dynamical exponent is $z=1$ for the Dirac QCP, and our framework can also determine $z$ when unknown (see SM Sec. S2 [53]), without any prior knowledge. We show the numerical results in Fig. 1b, in which the crossing point of $R_{\rm FM}$ versus $U$ for different $L$ precisely reveals a quantum phase transition from DSM to FM phase occurring at $U_{c}=7.220(37)$. This analysis fully aligns with the standard procedure of finite-size scaling, given the fixed relation $\tau=0.3L^{z}$. We then perform data collapse analysis of $R_{\rm FM}$ for different $L$ versus $(U-U_{c})L^{1/\nu}$ and achieve accurate critical exponent $1/\nu=1.18(3)$, as shown in Fig. 1c. Furthermore, fixing $U=U_{c}$, we calculate the imaginary-time evolution of FM structure order $S_{\rm FM}$ and fermion correlation $G_{\rm f}$ (defined in SM Sec. S2 [53]), whose scaling dimensions are $(1+\eta_{\phi})$ and $\eta_{\psi}$, respectively [56, 50]. By making the rescaled curves of $S_{\rm FM}L^{(1+\eta_{\phi})}$ and $G_{\rm f}L^{\eta_{\psi}}$ versus $\tau L^{-z}$ for different $L$ collapse according to Eq. (1) as shown in Figs. 1d and 1e, we achieve the anomalous dimensions of the order parameter boson $\eta_{\phi}=0.36(3)$ and the Dirac fermion $\eta_{\psi}=0.134(3)$, respectively.
The choice of the initial state and the imaginary-time length are two pivotal handles in our framework. Within the nonequilibrium scaling region near the QCP, the universal scaling form Eq. (1), as well as the resulting critical point and exponents, are independent of these choices, while the severity of the sign problem is sensitive to them. Flexible choices of the initial state and imaginary-time length not only help alleviate the sign problem to the greatest extent, but also provide a self-consistency check for the accuracy of the extracted critical point and exponents. In the demonstration shown in Figs. 1b and 1c, we used the DSM initial state with $\tau=0.3L^{z}$ to locate the critical point. The same result can be reproduced by changing $\tau$ to $\tau=0.5L^{z}$ or by switching the initial state to the FM state [53]. Moreover, the scaling collapses in Figs. 1d and 1e confirm that the system has entered the nonequilibrium scaling region. The details of self-consistency checks for different initial states and $\tau$ are provided in SM Sec. S2.2 [53]. Beyond the self-consistency checks, our estimates of the critical point and exponents are also approximately consistent with previous results from Gutzwiller QMC [56] and functional renormalization group [59]. The substantial mitigation of the sign problem enables us to simulate larger system sizes up to $L=23$, yielding reliable results (a systematic discussion is given in SM Sec. S1.5 [53]). Consequently, we unambiguously demonstrate that short-imaginary-time dynamics via QMC simulation offers a powerful approach to determine the phase transition point and critical exponents in sign-problematic strongly interacting models in a self-consistent, accurate, and unbiased manner, without relying on any prior knowledge.
Spinless $t$-$V$ model
@@
To further demonstrate the framework of our approach, we study another typical interacting Dirac-fermion model, termed as honeycomb spinless $t$-$V$ model, with the Hamiltonian:  [17, 60, 61, 62]:
@@
where $t$ is nearest-neighbor (NN) hopping amplitude and $V$ denotes the strength of NN density interaction. The appearance of the sign problem in Eq. (3) depends on the channel of Hubbard-Strotonovich (H-S) transformation. Hence, the model provides a genuine platform to confirm the accuracy and feasibility of our approach to unravel QCP in the presence of the sign problem. Previous sign-free QMC studies, with the H-S transformation in the hopping channel, reveal the critical properties of the QCP separating DSM and charge-density-wave (CDW) phases [61]. Here, we decouple the interaction in the sign-problematic density channel [63] and study the QCP via short-imaginary-time relaxation dynamics.
@@
Figure 3: Phase diagram and quantum criticality in $\rm SU(3)$ Hubbard model detected by short-imaginary-time dynamics with $\lambda_{8}$-AFM initial state. a, Phase diagram determined via short-imaginary-time dynamics. Insets show the energy spectra of DSM state (upper left) and sketch of $\lambda_{8}$-AFM order in which a fermion with one flavor (green “3”) is situated at one sublattice and double fermions with the other two flavors (pink “1” and violet “2”) are situated at the other sublattice (lower right). b-c, Critical point $U_{c}=1.10(5)$ for $\phi=0.075\pi$ and $1/\nu=0.68(5)$ determined via curves of $R_{\lambda_{8}\text{-AFM}}$ versus $U$ for different $L$ at $\tau=0.25L^{z}$, where $z=1$ and $\chi_{\nu}^{2}=1.239$ for the collapse quality. Shown in Inset of b is the evolution of average sign with red dashed curve marks $\tau=0.25L^{z}$. d-e, $\eta_{\phi}=0.55(5)$ and $\eta_{\psi}=0.15(3)$ determined via the scaling collapse of evolution curves of $S_{\lambda_{8}\text{-AFM}}$ and $G_{\rm f}$, respectively, with $\chi_{\nu}^{2}=1.382$ and $\chi_{\nu}^{2}=0.560$ demonstrating the collapse quality in the nonequilibrium scaling region to the right of the dashed lines.
@@
The procedure is the same as in the previous section, but here we demonstrate the case with the CDW initial state. The consistency check with the DSM initial state is provided in SM Sec. S3 [53]. We fix $\tau=0.3L^{z}$, and determine QCP from the crossing point of curves of correlation-length ratio $R_{\rm CDW}$ versus $V$ ($R_{\rm CDW}$ is defined in SM Sec. S3 [53]), as displayed in Fig. 2a, giving rise to $V_{c}=1.35(1)$. The scaling collapse analysis determines the value of $\nu$ as $1/\nu=1.30(18)$. Additionally, the anomalous dimensions $\eta_{\phi}=0.49(5)$ and $\eta_{\psi}=0.073(4)$ are obtained from the evolution of the CDW structure factor $S_{\rm CDW}$ and the fermion correlation $G_{\rm f}$ (defined in SM Sec. S3 [53]), respectively, at $V_{c}$. Both $V_{c}$ and $\eta_{\phi}$ are consistent with previous numerical results [60, 61, 59]. Notably, $\eta_{\psi}$ is numerically determined for the first time, aligning consistently with the previous results of the functional renormalization group [59]. These results for model (3) further establish the feasibility of the approach based on short-imaginary-time dynamics in accessing the critical properties, even in the presence of the sign problem. The ability to select different initial states also offers valuable benchmarks, thereby further solidifying the reliability of our approach.
SU(3) Hubbard model
@@
Building on the successful application of our method to previous models, we now apply it to systematically investigate the unexplored sign-problematic systems. Our particular interest lies in identifying an exotic Dirac QCP belonging to a novel universality class. To this end, we consider $\rm SU(3)$ Hubbard model on square lattice with staggered magnetic flux, described by the following Hamiltonian: